how do you know if a quadratic equation will have one, two or no solutions? If only given the solution, how do you find a quadratic equation? Is it possible to have different quadratic equations with the same solution?
Answer by stanbon(75887)

how do you know if a quadratic equation will have one, two or no solutions?
Check the determinant: b^2-4ac
--------------------------------
If only given the solution, how do you find a quadratic equation?
If a is a solution, (x-a) is a factor
---------------------------------------------
Is it possible to have different quadratic equations with the same solution?
Ans: Yes; y = (x-2)(x+3) and y = 2(x-2)(x+3) are two different equations
that have the same solutions.
=======================================
Cheers,
Stan H.